Please help me solve this problem "Average age of 4 men, 4 years ago was 49 years. After 4 year a man of age 42 join them. Now what is average age?"
Answer by Cromlix(4381)

Average age of 4 men, 4 years ago was 49 years.
Total ages = 4 * 49 = 196 years
After 4 years Total ages = 196 + 4 = 200
A man aged 42 joins them 200 + 42
= 242
Average = 242/5 = 48.4